find f+g, for f(x)=3x-6 and g(x)=4x+1
Found 2 solutions by JBarnum, drk:
Answer by JBarnum(2146)

find f+g, for f(x)=3x-6 and g(x)=4x+1
this is simple substitution
f+g
(3x-6)+(4x+1)
(7x-5) is the answer


to check this answer simply make up a value for x
if x=1
f=3(1)-6=3-6=-3
g=4(1)+1=4+1=5
f+g=5+(-3)=2
if (7x-5)=f+g
then(7(1)-5)=7-5=2
2=2
correct


Answer by drk(1908)

we add the functions to get
3x - 6 + 4x + 1 = 7x - 5